OpenStudy (anonymous):

6. If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

Substitute in 5 for x and check the math.

OpenStudy (anonymous):

why 5?

OpenStudy (anonymous):

5*5 + 2(5)-35=0 25+10-35=0 35-35=0 0=0

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[\qquad\downarrow\]\[x_{1,2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]